OpenStudy (anonymous):

The product of twice a number and three is the same as the difference of five times the number and 5/7. Find the number

OpenStudy (anonymous):

Do you have a problem with finding the equation for the number, or with solving the equation itself?

OpenStudy (anonymous):

finding the equation for the number i think i have it but i'm not really sure

OpenStudy (anonymous):

What do you have so far?

OpenStudy (anonymous):

2x+3=5x+\[\frac{ 5 }{ 7}\]

OpenStudy (anonymous):

i feel like i'm wrong though kinda stuck

OpenStudy (anonymous):

Let's go over it then :) "The product of twice a number and three is the same as the difference of five times the number and 5/7." The product of twice a number and three: \[2x * 3\] is the same as (=) the difference of 5 times the number and 5/7: \[5x - \frac{5}{7}\] So \[2x * 3 = 5x - \frac{5}{7}\]

OpenStudy (anonymous):

oh ok i got it now thanks! :)